Albumin is being separated from Immunoglobin G (IgG) in a chromatography column
having a stationary phase fraction of 0.25. The mobile phase retention time is 10
min. The distribution coefficients for albumin and IgG are 0.1 and 1, respectively. If
the albumin peak has a characteristic width of 0.52 minutes, predict the selectivity
and resolution.
1
Expert's answer
2020-12-29T02:42:59-0500

the volume of the column is 39ml calculated using the given data

the mobile phase retention time can be calculated using equation

tM = "\frac{V_Cε}{Q}" "\implies" "\frac{39(0.25)}{10}" = 0.975 min

the retention times can be calculated using equation

tR = tM {1+ ( "\frac{1-ε}{ε}" ) K}


tR,alb = 0.975 { 1 + ("\frac{1-0.25}{0.25}" ) 0.1 } = 1.25 min


tR,IgG = 0.975 { 1 + ("\frac{1-0.25}{0.25}") 1 } = 3.9 min


the number of theoretical plates in the column can be calculated from albumin retention time data using equation

N = 16("\frac{t_g}{w}" )2 "\implies" 16 ("\frac{1.27}{0.52}")2 = 95


the peak width of IgG can be calculated using equation

WIgG = "\frac{t_{R.IgG}}{\sqrt\frac{N}{16}}" "\implies" "\frac{3.9}{\sqrt\frac{95}{16}}" = 1.6 min


the resolution of separation can be calculated using equation

R = "\frac{t_{R2}-t_{g1}}{0.5(w_1+w_2)}" "\implies" "\frac{3.9-1.27}{0.5(1.6+0.52)}" = 2.58


the selectivity can be calculated using equation

α = "\frac{K_2}{K_1}" = "\frac{t_{R2}-t_M}{t_{R1}-t_M}" "\implies" "\frac{1}{0.1}" = 10


the height of the theoritical plate can be calculated using equation

H = "\frac{l}{N}" = "\frac{50}{95}" = 0.526 cm